# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

### v1

@@ Abstract
Despite the increasing use of large language models for creative tasks, their outputs often lack diversity. Common solutions, such as sampling at higher temperatures, can compromise the quality of the results. Drawing on information theory, we propose a context-based score to quantitatively evaluate value and originality. This score incentivizes accuracy and adherence to the request while fostering divergence from the learned distribution. We propose using our score as a reward in a reinforcement learning framework to fine-tune large language models for maximum performance. We validate our strategy through experiments in poetry generation and math problem solving, demonstrating that it enhances the value and originality of the generated solutions.
Thinking Outside the (Gray) Box: A Context-Based Score for
Assessing Value and Originality in Neural Text Generation
Giorgio Franceschelli University of Bologna giorgio.franceschelli@unibo.it                        Mirco Musolesi University College London University of Bologna m.musolesi@ucl.ac.uk
1 Introduction
@@
Foundation models Bommasani et al. (2021), particularly large language models (LLMs) Bubeck et al. (2023); Gemini Team et al. (2023); Touvron et al. (2023), are significantly transforming creative activities. They can serve as a foundation for co-creation systems involving human and artificial authors Lin et al. (2023); they can be utilized to generate software code Rozière et al. (2023); they can even be employed to foster scientific research Boiko et al. (2023). However, the nature of the self-supervised learning algorithms used for the training of these models tend to make them to generate samples as close as possible to the training data distribution Franceschelli and Musolesi (2024a). In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) Christiano et al. (2017), is often necessary to generate appropriate and accurate responses. However, this process tends to further reduce output diversity Kirk et al. (2024), and linguistic creativity tends to be lower than that of humans Lu et al. (2025). On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request. One typical solution is to sample at higher temperature to increase diversity. However, this might lead to generating incoherent text Peeperkorn et al. (2024).
@@
In order to address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs. The definition of CoVO is grounded in the analysis of mutual information MacKay (2003) between the model’s outputs and inputs, and vice versa. More specifically, we formulate a new optimization problem where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model. In this way, we optimize for solutions that are appropriate for the input request but also different from the outputs we would normally obtain from the model. In particular, we show that our information-theoretic score can be used as a reward in RL-based fine-tuning algorithms, guiding pre-trained models toward more diverse yet valuable solutions.
We also present the theoretical foundations of our approach and we discuss several strategies for the practical implementation of the proposed approach and its adaptation to current LLMs. Experiments on math problem solving and poetry generation show that our score correlates with domain-specific measures for value and originality, and that our approach can increase the quality and diversity of outputs, presenting itself as a candidate for creativity applications of current foundation models.
The remainder of the paper is structured as follows. First, we provide an overview of the state of the art in this area (Section 2) and introduce the key concepts and background of our work (Section 3). Then, we discuss the definition of CoVO (Section 4) and show different ways to use it in practice as a reward for fine-tuning autoregressive models based on RL (Section 5). We validate the proposed solutions on poetry generation and math problem solving (Section 6). Finally, we discuss future work and limitations of our approach (Sections 7 and 8).
2 Related Work
2.1 Information Theory and Creativity
@@
The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades. Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi, 2024a). However, these methods are often domain-specific and typically require substantial human effort to implement and assess. In contrast, solutions based on information theory Shannon (1948); Cover (1999) offer a more universally applicable approach.
@@
Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains. Bayesian surprise Baldi and Itti (2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty França et al. (2016); Varshney et al. (2019) or surprise Mazzaglia et al. (2022); Schmidhuber (2010). Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures. Surprisal Tribus (1961), i.e., Shannon’s self-information, has also been used Bunescu and Uduehi (2019); Fernandez Monsalve et al. (2012); Barto et al. (2013) extensively discuss surprisal and Bayesian surprise, and how novelty differ from them. Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more Achiam and Sastry (2017). Sun et al. (2025) apply this idea to improve exploration in RLHF Christiano et al. (2017). Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience. Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness Li et al. (2016). However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
2.2 LLMs and Creativity
@@
Since the introduction of GPT models Brown et al. (2020); OpenAI (2023) and their competitors (e.g., Touvron et al., 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this Franceschelli and Musolesi (2024b). For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs Stevenson et al. (2022) and to investigate methods for enhancing their performance Goes et al. (2023); Summers-Stay et al. (2023). Porter and Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones. In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal. Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems Popescu-Belis et al. (2023) or utilized in zero-shot settings to emulate the writing styles of famous authors Sawicki et al. (2023). It has also been shown that these models can be fine-tuned via RLHF Christiano et al. (2017) to write short poems that human evaluators find more creative Pardinas et al. (2023). Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human Ding et al. (2023) or AI Bradley et al. (2024) feedback to measure the quality of the generated outputs.
3 Preliminaries
3.1 Language Modeling
@@
3.2 Reinforcement Learning for Language Models
@@
Due to its adherence to the formal framework of Markov decision processes Sutton and Barto (2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation Bachman and Precup (2015). The LLM plays the role of the agent, and each generated token represents an action $a_{t}$. The current version of the generated output $\mathbf{x_{t}}$ is part of the state $\mathbf{s_{t}}$ (potentially with additional information such as initial prompts). Finally, the reward $r_{t+1}$ measures the “quality” of the current output. A common strategy is to assign a zero reward for each $\mathbf{x}_{t},t\neq T$ and a sentence-based reward when the final output is generated. Within this framework, any policy-based method, such as Proximal Policy Optimization (PPO) Schulman et al. (2017), can be employed to train or fine-tune the LLM to optimize a given objective. Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences Franceschelli and Musolesi (2024c). To avoid deviating too much from the pre-trained model while also encouraging exploration, the full reward function usually considers a Kullback-Leibler (KL) divergence term as follows Stiennon et al. (2020):
$R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x}|\mathbf{z})},$ (1)
where $\boldsymbol{\theta}^{PT}$ is the set of parameters of the pre-trained model, $\beta$ weighs the KL penalty, and $r(\mathbf{z},\mathbf{x})$ can be any non-differentiable function to score the model’s output.
In particular, RLHF Christiano et al. (2017) has been widely adopted as the final training stage in various popular LLMs to align them with human preferences, i.e., by using a learned reward function that approximates human feedback. While highly effective, RLHF suffers from several open problems Casper et al. (2023). These include obtaining accurate, unbiased, and representative human feedback, as well as making the process resource-efficient. Moreover, RLHF tends to be complex and unstable due to the additional requirement of training a reward model and its reliance on RL. To address these challenges, Direct Preference Optimization (DPO) Rafailov et al. (2023) directly optimizes the language model without the need for explicit reward modeling or RL. DPO implicitly performs reward maximization with a KL-divergence penalty through the following loss:
$\mathcal{L}_{DPO}=\min(-\mathbb{E}_{(\mathbf{z},\mathbf{x_{w}},\mathbf{x_{l}})\sim X}[\log\sigma($ $\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x_{w}}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x_{w}}|\mathbf{z})}$ (2)
-- $\beta\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x_{l}}|\mathbf{z})}{p_{\boldsymbol{\theta}^{PT}}(\mathbf{x_{l}}|\mathbf{z})})])$
where $\sigma$ is the sigmoid function, $\beta$ is a parameter controlling the deviation from the base model, and $\mathbf{x_{w}},\mathbf{x_{l}}$ are the chosen and rejected outputs, respectively. In other words, DPO fits an implicit reward whose optimal policy is the language model itself.
4 A Context-Based Score for Valuable and Original Generation
@@
where $I(\mathbf{x},\mathbf{y})$ is just $I(\mathbf{x},\mathbf{y},1,1)$. By applying the Bayes theorem, i.e., $\log p(a|b)=\log p(b|a)+\log p(a)-\log p(b)$, we can substitute the $\log p(\mathbf{y})$ term as follows:
@@
Let us now consider the case where $\lambda_{v},\lambda_{o}>0$, for example, $\lambda_{v}=\lambda_{o}=1$. Solving this maximization problem involves finding the target $\mathbf{y}$ that maximizes the posterior probability of $\mathbf{x}$ while also being unlikely given $\mathbf{x}$. In other words, the optimal $\mathbf{y}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$. While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal Tribus (1961), is widely used to measure diversity and surprise Barto et al. (2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness. Indeed, if the request (e.g., a problem or a task) can be easily predicted from the outcome, the outcome must be a (good) example of that task or a correct solution for that problem. While other prominent definitions such as Boden’s Boden (2003) seek to include a third requirement for creativity, i.e., novelty, we believe that by considering only a specific context, novelty and surprise become indistinguishable. In particular, being novel in a specific context $\mathbf{x}$ means doing something different from what has been previously experienced under $\mathbf{x}$ by the creator. However, since a self-supervised-trained model can be seen as a compression of training data Franceschelli et al. (2024), $-\log P(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty. Therefore, we choose to refer to this term as originality, as it encompasses both novelty and surprise. On the other hand, $\log P(S|T)$ is less ambiguous and maps directly into value.
We now provide a formal definition of our proposed score. In summary, the CoVO (Context-based Value and Originality) score for a target $\mathbf{y}$ given a source $\mathbf{x}$ on a reference probability distribution $p$ is defined as:
CoVO Score $s_{CoVO}(\mathbf{x},\mathbf{y},p)=\underbrace{\lambda_{v}\log p(\mathbf{x}|\mathbf{y})}_{\textrm{Value}}\underbrace{-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})}_{\textrm{Originality}}$ (9)
5 Implementation and Optimization with Autoregressive Models
We now discuss the implementation of the CoVO score with autoregressive models. Using the notation introduced in Section 3.1, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$. However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences. To avoid this, we propose to use the $T$-th root: $T]{\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$. By leveraging the properties of the logarithm, we obtain:
@@ 5 Implementation and Optimization with Autoregressive Models
It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\boldsymbol{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$. In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity. Since $\mathbf{y}$ has been sampled from $p_{\boldsymbol{\theta}}$, its probability would be high by definition. However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
@@
Inspired by Macedo et al. (2004), we propose to normalize $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$. For probabilities, $n_{min}=0$, while $n_{max}=\max p_{\boldsymbol{\theta}}(b)$, thus obtaining the overall mapping for $p_{\boldsymbol{\theta}}$: $\frac{p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max p_{\boldsymbol{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$. Once again, by applying the properties of logarithms, we obtain:
CoVO Score with Autoregressive Models $s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11) $\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$ $\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$
Despite these adjustments, the two components of our score still exhibit fundamental differences in the magnitude of variations. For instance, given the same source $\mathbf{x}$, small variations in the target $\mathbf{y}$ result in larger fluctuations in the originality component compared to the value component. To address this issue, we propose to normalize the two parts independently. We implement and test two distinct methodologies: single-batch normalization, which standardizes each component using its mean and standard deviation within a single batch; and full-training normalization, which standardizes each component using the mean and standard deviation calculated over the entire training period, maintaining running statistics across batches.

### v2

@@ Abstract
Despite the increasing use of large language models for creative tasks, their outputs often lack diversity. Common solutions, such as sampling at higher temperatures, can compromise the quality of the results. Dealing with this trade-off is still an open challenge in designing AI systems for creativity. Drawing on information theory, we propose a context-based score to quantitatively evaluate value and originality. This score incentivizes accuracy and adherence to the request while fostering divergence from the learned distribution. We show that our score can be used as a reward in a reinforcement learning framework to fine-tune large language models for maximum performance. We validate our strategy through experiments considering a variety of creative tasks, such as poetry generation and math problem solving, demonstrating that it enhances the value and originality of the generated solutions.
Introduction
Figure 1: A summary of our method: the target model (orange) produces $G$ outputs for each prompt; a frozen reference model (blue) computes the value and originality for each output; the overall scores are used in GRPO to correct the target model (orange line).
@@ Introduction
Foundation models (Bommasani et al. 2021), particularly large language models (LLMs) (Bubeck et al. 2023; Gemini Team et al. 2023; Touvron et al. 2023), are significantly transforming creative activities. They can serve as a foundation for co-creation systems involving human and artificial authors (Lin et al. 2023); they can be utilized to generate software code (Rozière et al. 2023); they can even be employed to foster scientific research (Boiko, MacKnight, and Gomes 2023). However, the nature of the self-supervised learning algorithms used for the training of these models tends to make them generate samples as close as possible to the training data distribution (Franceschelli and Musolesi 2024a). In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) (Christiano et al. 2017), is often necessary to generate appropriate and accurate responses. However, this process tends to further reduce output diversity (Kirk et al. 2024), and linguistic creativity tends to be lower than that of humans (Lu et al. 2025). On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request. One typical solution is to sample at a higher temperature to increase diversity. However, this might lead to generating incoherent text (Peeperkorn et al. 2024).
@@
In order to address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs. The definition of CoVO is grounded in the analysis of mutual information (MacKay 2003) between the model’s outputs and inputs, and vice versa. More specifically, we formulate a new optimization problem where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model. In this way, we optimize for solutions that are appropriate for the input request but also different from the outputs we would normally obtain from the model. In particular, we show that our information-theoretic score can be used as a reward in RL-based fine-tuning algorithms, guiding pre-trained models toward more diverse yet valuable solutions. Figure 1 summarizes our proposed approach.
In summary, our key contributions are the following:
  • We present the theoretical foundations of our approach, deriving our context-based score for value and originality from the concept of mutual information.
  • We discuss how the score can be practically computed in the case of autoregressive models, and how it can be used as a reward with Group Relative Policy Optimization (GRPO) (Shao et al. 2024), a state-of-the-art reinforcement learning algorithm for fine-tuning LLMs.
  • We evaluate our GRPO-based method on mathematical problem solving, poetry generation, and the tasks included in NoveltyBench (Zhang et al. 2025), demonstrating that our scoring metric correlates with domain-specific measures of value and originality. Moreover, we show that our approach enhances both the quality and diversity of generated outputs, positioning it as a strong candidate for creativity-focused applications of current foundation models.
Related Work
Information Theory and Creativity
@@
The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades. Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi 2024a). However, these methods are often domain-specific and typically require substantial human effort to implement and assess. In contrast, solutions based on information theory (Shannon 1948; Cover 1999) offer a more universally applicable approach.
@@
Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains. Bayesian surprise (Baldi and Itti 2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty (França et al. 2016; Varshney et al. 2019) or surprise (Mazzaglia et al. 2022; Schmidhuber 2010). Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures. Surprisal (Tribus 1961), i.e., Shannon’s self-information, has also been used (Bunescu and Uduehi 2019; Fernandez Monsalve, Frank, and Vigliocco 2012); Barto, Mirolli, and Baldassarre (2013) extensively discuss surprisal and Bayesian surprise, and how novelty differs from them. Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more (Achiam and Sastry 2016). Sun et al. (2025) apply this idea to improve exploration in RLHF (Christiano et al. 2017). Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience. Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness (Li et al. 2016). However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
LLMs and Creativity
@@
Since the introduction of GPT models (Brown et al. 2020; OpenAI 2023) and their competitors (e.g., Touvron et al. 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this (Franceschelli and Musolesi 2024b). For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs (Stevenson et al. 2022) and to investigate methods for enhancing their performance (Goes et al. 2023; Summers-Stay, Voss, and Lukin 2023). Porter and Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones. In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal. Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems (Popescu-Belis et al. 2023) or utilized in zero-shot settings to emulate the writing styles of famous authors (Sawicki et al. 2023). It has also been shown that these models can be fine-tuned via RLHF (Christiano et al. 2017) to write short poems that human evaluators find more creative (Pardinas et al. 2023). Reinforcement learning (RL) can also encourage large language models (LLMs) to produce more diverse outputs (Chung et al. 2025), and can be leveraged to optimize metrics commonly linked to creative expression (Ismayilzada et al. 2025). Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human (Ding et al. 2023) or AI (Bradley et al. 2024) feedback to measure the quality of the generated outputs.
Preliminaries
Language Modeling
@@
Reinforcement Learning for Language Models
@@
Due to its adherence to the formal framework of Markov decision processes (Sutton and Barto 2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation (Bachman and Precup 2015). The LLM plays the role of the agent, and each generated token represents an action $a_{t}$. The current version of the generated output $\mathbf{x_{t}}$ is part of the state $\mathbf{s_{t}}$ (potentially with additional information such as initial prompts). Finally, the reward $r_{t+1}$ measures the “quality” of the current output. A common strategy is to assign a zero reward for each $\mathbf{x}_{t},t\neq T$ and a sentence-based reward when the final output is generated. Within this framework, any policy-based method can be employed to train or fine-tune the LLM to optimize a given objective. Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences (Franceschelli and Musolesi 2024c).
A widely used RL algorithm for fine-tuning LLMs is Proximal Policy Optimization (PPO) (Schulman et al. 2017), which aims to maximize the following objective:
$\mathcal{J}(\boldsymbol{\theta})=\mathbb{E}_{t}\!\left[\min(r_{t}(\boldsymbol{\theta})\hat{A}_{t},\mathtt{clip}(r_{t}(\boldsymbol{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$ (1)
where $r_{t}(\boldsymbol{\theta})=\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}{\pi_{\boldsymbol{\theta}_{old}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}$ with $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}_{old}}$ denoting the current and old policy models, repectively; $\mathbf{x}$ is the output sampled from the old policy given the prompt $\mathbf{z}$; and $\epsilon$ is a clipping factor used to stabilize training. The advantage $\hat{A}_{t}$ is usually computed through Generalized Advantage Estimation (Schulman et al. 2016) based on the full rewards $R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{\pi_{ref}(\mathbf{x}|\mathbf{z})}$, thus integrating a KL penalty with respect to a reference model (usually, the same model before fine-tuning), and a learned value function $v_{\boldsymbol{\phi}}(\mathbf{s})$. However, learning such a value function is computationally intensive, and its training is complicated by the fact that only the last state is scored by the reward function. Moreover, the inclusion of the KL penalty as an auxiliary reward term adds complexity to the advantage estimation process. To address these issues, Group Relative Policy Optimization (GRPO) (Shao et al. 2024) has been introduced. GRPO directly adds a KL divergence term $-\beta D_{KL}(\pi_{\boldsymbol{\theta}}||\pi_{ref})$ to the loss rather than to the single rewards, and especially obviates the need for a value function approximator by using the average reward of multiple sampled outputs as the baseline:
$\hat{A}_{i,t}=\frac{r_{i}-\mathtt{mean}(\mathbf{r})}{\mathtt{std}(\mathbf{r})},$ (2)
with $\mathbf{r}=\{r_{1},r_{2},...,r_{G}\}$ as the list of rewards received by each of the $G$ outputs sampled from the same prompt.
A Context-Based Score for Valuable and Original Generation
@@
where $I(\mathbf{x},\mathbf{y})$ is just $I(\mathbf{x},\mathbf{y},1,1)$. Computing the absolute probability $p(\mathbf{y})$ can be difficult, as usually generative models are developed to assign conditional probabilities. By applying the Bayes theorem, i.e., $\log p(a|b)=\log p(b|a)+\log p(a)-\log p(b)$, we can substitute the $\log p(\mathbf{y})$ term as follows:
@@
Let us now consider the case where $\lambda_{v},\lambda_{o}>0$, for example, $\lambda_{v}=\lambda_{o}=1$. Solving this maximization problem involves finding the target $\mathbf{y}$ that maximizes the posterior probability of $\mathbf{x}$ while also being unlikely given $\mathbf{x}$. In other words, the optimal $\mathbf{y}^{*}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$. While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal (Tribus 1961), is widely used to measure diversity and surprise (Barto, Mirolli, and Baldassarre 2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness. Indeed, if the request (e.g., a problem or a task) can be easily predicted from the outcome, the outcome must be a (good) example of that task or a correct solution for that problem. While other prominent definitions, such as Boden’s, seek to include a third requirement for creativity, i.e., novelty (Boden 2003), we believe that by considering only a specific context, novelty and surprise become indistinguishable. In particular, being novel in a specific context $\mathbf{x}$ means doing something different from what has been previously experienced under $\mathbf{x}$ by the creator. However, since a self-supervised-trained model can be seen as a compression of training data (Franceschelli, Cevenini, and Musolesi 2024), $-\log p(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty. Therefore, we choose to refer to this term as originality, as it encompasses both novelty and surprise. On the other hand, $\log p(\mathbf{x}|\mathbf{y})$ is less ambiguous and maps directly into value.
In summary, the CoVO (Context-based Value and Originality) score for a target $\mathbf{y}$ given a source $\mathbf{x}$ on a reference probability distribution $p$ is formally defined as:
$s_{CoVO}(\mathbf{x},\mathbf{y},p)=\underbrace{\lambda_{v}\log p(\mathbf{x}|\mathbf{y})}_{\textrm{Value}}\underbrace{-\lambda_{o}\log p(\mathbf{y}|\mathbf{x})}_{\textrm{Originality}}$ (9)
Implementation and Optimization with Autoregressive Models
We now discuss the implementation of the CoVO score with autoregressive models. Using the notation introduced above, in the context of a $\boldsymbol{\theta}$-parameterized LLM, $p(\mathbf{y}|\mathbf{x})$ can be expressed as $\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})$. However, considering just the product of all the conditioned probabilities for an optimization problem would lead to preferring shorter sequences. To avoid this, we propose to use the $T$-th root: $T]{\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}$. By leveraging the properties of the logarithm, we obtain:
@@
It is worth noting that the vocabulary of an LLM can be extremely large, which can cause $p_{\boldsymbol{\theta}}(a|b)$ to be small even when $a$ is the most probable event given $b$. In particular, when an LLM generates $\mathbf{y}$ given $\mathbf{x}$ and then evaluates both $p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$ and $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$, this can lead to a significant discrepancy between the magnitude of value and diversity. Since $\mathbf{y}$ has been sampled from a model based on $p_{\boldsymbol{\theta}}$, its probability would be high by definition. However, there may be various ways (possibly through synonyms) to define $\mathbf{y}$, leading to a smaller probability of $\mathbf{x}$.
@@
Inspired by Macedo, Reisenzein, and Cardoso (2004), we propose to counteract this problem by normalizing $p_{\boldsymbol{\theta}}(a|b)$ via $n^{\prime}=\frac{n-n_{min}}{n_{max}-n_{min}}$. For probabilities, $n_{min}=0$, while $n_{max}=\max p_{\boldsymbol{\theta}}(b)$, thus obtaining the overall mapping for $p_{\boldsymbol{\theta}}$: $\frac{p_{\boldsymbol{\theta}}(y_{t}|\mathbf{y}_{<t},\mathbf{x})}{\max p_{\boldsymbol{\theta}}(\mathbf{y}_{<t},\mathbf{x})}$. Once again, by applying the properties of logarithms, we obtain:
$s_{CoVO}^{AR_{norm}}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=\lambda_{v}s_{v}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})+\lambda_{o}s_{o}(\mathbf{x},\mathbf{y},p_{\boldsymbol{\theta}})=$ (11)
$\lambda_{v}\frac{\sum_{i=1}^{|\mathbf{x}|}(\log p_{\boldsymbol{\theta}}(x_{i}|\mathbf{x}_{<i},\mathbf{y})-\max\log p_{\boldsymbol{\theta}}(\mathbf{x}_{<i},\mathbf{y}))}{|\mathbf{x}|}-$
$\lambda_{o}\frac{\sum_{j=1}^{|\mathbf{y}|}(\log p_{\boldsymbol{\theta}}(y_{j}|\mathbf{y}_{<j},\mathbf{x})-\max\log p_{\boldsymbol{\theta}}(\mathbf{y}_{<j},\mathbf{x}))}{|\mathbf{y}|}.$